OpenStudy (anonymous):

Help me please! I want to learn how to do this! What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (anonymous):

@dmezzullo @reemii ??

OpenStudy (reemii):

parallel lines have same slope. what is the slope of the second line?

OpenStudy (anonymous):

I'm not sure :/

OpenStudy (anonymous):

firstly you must get the gradient of the first line which is 5/2...since we know that when two lines are parallel their slopes are equal...the slope of the second line is also 5/2

OpenStudy (anonymous):

5/3 i mean

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (reemii):

ehm. \(\frac{6-(-4)}{4-(-1)} = \frac{10}{5} = 2\). isn't the slope =2 ?

OpenStudy (anonymous):

Is it?

OpenStudy (reemii):

your opinion? what is \(\Delta y\)? what is \(\Delta x\)

OpenStudy (anonymous):

Yeah I agree with you! Another question?

OpenStudy (anonymous):

dude firstly do you understand the concept

OpenStudy (anonymous):

Yes I do, now I do

OpenStudy (anonymous):

cool dude....cheers to that..wink..